Suppose we deposit $565 worth of currency into the bank. Banks make as many loans as possible, people hold no additional currency, and the reserve ratio is 15%. As a result of the loan process, what is the value of total deposits in the banking system? Enter a whole number with no dollar sign. Round to the nearest whole number.
1
Expert's answer
2020-11-24T06:23:06-0500

total deposit = initial deposit "\times" money multiplier

"=565 \times 1/ reserve" "requirement"

"=565 \times\frac{1}{0.15} = 3766.67"

"=3767"